OpenStudy (anonymous):

Which of the following points are solutions to the system of inequalities shown below? Check all that apply. y≤-2x+6 x>1 A) (1,4) B) (2,11) C) (2,1) D) (0,6) E) (3,0) F) (5,-6)

OpenStudy (radar):

You will need to examine each one. A. x is not greater than 1 so I would believe it can be disregarded. The same goes for D.

OpenStudy (radar):

Looking at B. It satisfies the x>1 now for the other test 11< or = -8+6 or\[11\le 6\]It fails.

OpenStudy (radar):

Check out C now \[1\le -2\times2+6\]\[1\le 2\] That is a good one (a solution). Hopefully you now get the drift and can check out E and F

OpenStudy (anonymous):

Thank you sir. :)

OpenStudy (radar):

No problem, if you run into difficulty, I will be standing by.

OpenStudy (anonymous):

So does E equal the inequality? And F does not?

OpenStudy (radar):

let me check

OpenStudy (radar):

\[0\le -2\times 3 +6\] 0=0 it is ok and is a solution. \[-6\le -2\times 5 +6\]\[-6\le -4\] That is also true thus a good answer also.

OpenStudy (radar):

So E and F are OK and are solutions also.

OpenStudy (anonymous):

Thank you!!! :D